# Scaling Laws for Uncertainty in Deep Learning

### v1

@@ A.2.1 Asymptotic behaviors
where $\lambda$ is a positive rational number, $m$ a positive integer and $\mathbb{E}_{n}[\cdot]$ is the expectation operator the overall dataset. The constant $\lambda$ is called a learning coefficient since it is dominant in the leading terms of the equations above, which represents the $\mathbb{E}_{n}[G_{n}]-n$ and $F_{n}-n$ learning curves. $m$ is called a multiplicity in algebraic geometry. In the case of regular models $\lambda=\frac{d}{2}$ and $m=1$ where $d$ is the number of model parameters.
Appendix B Additional results
@@
Figure 9: Vision Transformer (ViT) on ImageNet-32 dataset: We use mc dropout with fixed dropout rate $p=0.1$. We train for $200$ epochs using SGD with cosine annealing. In (a) we report uncertainty scaling on subsets from size $250$K up to $1.2$M. In (b) we discard the first point highlighting the decreasing rate of eu for large values of $N$.
Figure 10: ResNets on CIFAR-100 dataset: We train the models with mc dropout with fixed dropout rate $p=0.2$ (first row) and $p=0.5$ (second row). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data. We report results from $10$ independent folds (varying both data subsampling and model initialization) showing the mean uncertainty for each $N$ subset. The dashed lines represent linear regressions fitted to the average (over the folds for every $N$) of each uncertainty metric. Both axes are on a logarithmic scale.
@@
Figure 11: WideResNets on CIFAR-100 dataset: We train the models with mc dropout with fixed dropout rate $p=0.2$ (first row) and $p=0.5$ (second row). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data and we show the mean uncertainty for each $N$ subset. Dashed lines represent linear regression fitted to the test mean of each uncertainty metric. Both axes are on a logarithmic scale.
@@
Figure 12: WideResNets on CIFAR-10 dataset - extrapolation: We report uncertainty trends for different WideResNets and dropout rates. We mention in the main paper that the scaling laws we derive in this context are practically useful to extrapolate uncertainties to $N$ arbitrarily large. We report $\mathbb{V}[\sigma(\mathbf{f})]$, the variance computed over the Softmax predictions of the mc samples, averaged over all the test samples ($\times$ in the plots). The black dashed lines have slope $\gamma_{f}$, the blue dashed lines correspond to two thresholds of predictive uncertainty.
@@
Figure 13: ResNets on CIFAR-100 dataset with Deep Ensembles: We train the models with Deep Ensembles with ensemble members $M=5$ (first row) and $M=10$ (second row). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data and we show the mean uncertainty for each $N$ subset. Dashed lines represent linear regression fitted to the test mean of each uncertainty metric. Both axes are shown on a logarithmic scale.
@@
Appendix C Experimental setup
Table 1: Vision Transformer (ViT) architecture - implementation from https://github.com/kentaroy47/vision-transformers-cifar10.

### v2

@@ A.2.1 Asymptotic behaviors
where $\lambda$ is a positive rational number, $m$ a positive integer and $\mathbb{E}_{n}[\cdot]$ is the expectation operator the overall dataset. The constant $\lambda$ is called the rlct or learning coefficient, since it is dominant in the leading terms of the equations above, which represents the $\mathbb{E}_{n}[G_{n}]-n$ and $F_{n}-n$ learning curves. In algebraic geometry, $m$ is called a multiplicity. In the case of regular models $\lambda=\frac{d}{2}$ and $m=1$, where $d$ is the number of model parameters.
A.2.2 On connecting slt to uncertainty scaling in deep models
Our use of Bayesian linear regression is primarily to illustrate how, in a simplified setting, generalization error can be cleanly decomposed into aleatoric and epistemic components, thereby exposing uncertainty measures that are otherwise less apparent. We are aware that this only scrapes the surface of the problem, as extending these insights to deep, over-parameterized networks remains an open challenge. Within the broader perspective of slt (Watanabe, 2009), neural networks form a singular and non-realizable class of models, for which theoretical results—such as asymptotic error bounds or explicit learning coefficients—are scarce. In particular, Bayes and Gibbs errors are governed by the rlct, which quantifies effective dimensionality; yet for over-parameterized architectures, the roles of rlct and related invariants remain poorly understood despite recent advances in local approximations, such as the Local Learning Coefficient (llc) (Lau et al., 2024).
Appendix B Uncertainty scaling: a study on linearized Laplace
B.1 Setup
We consider a classification problem with $C$ classes on a dataset $\mathcal{D}=\{({{\boldsymbol{\mathbf{x}}}}_{n},y_{n})\}_{n=1}^{N}$, where inputs ${{\boldsymbol{\mathbf{x}}}}_{n}\in\mathbb{R}^{D}$ and targets $y_{n}\in\{1,\dots,C\}$. The model is a Bayesian neural network (bnn) with parameters ${{\boldsymbol{\mathbf{\theta}}}}\in\mathbb{R}^{P}$ mapping inputs to a vector of logits ${{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}})\in\mathbb{R}^{C}$. The likelihood is defined as $p(\mathcal{D}\mid{{\boldsymbol{\mathbf{\theta}}}})=\prod_{n=1}^{N}p(y_{n}\mid{{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}}_{n};{{\boldsymbol{\mathbf{\theta}}}}))$. The log-joint density (unnormalized posterior) can be written as:
$\ell({{\boldsymbol{\mathbf{\theta}}}},\mathcal{D})=\sum_{n=1}^{N}\log p(y_{n}\mid{{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}}_{n};{{\boldsymbol{\mathbf{\theta}}}}))+\log p({{\boldsymbol{\mathbf{\theta}}}}).$ (41)
The Laplace approximation constructs a Gaussian surrogate $q({{\boldsymbol{\mathbf{\theta}}}})\approx p({{\boldsymbol{\mathbf{\theta}}}}\mid\mathcal{D})$ by matching the curvature of the true posterior at its peak. This is achieved via a second-order Taylor expansion of $\ell({{\boldsymbol{\mathbf{\theta}}}},\mathcal{D})$ around the maximum a posteriori (Maximum a posteriori (map)) estimate ${{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}$:
$\ell({{\boldsymbol{\mathbf{\theta}}}},\mathcal{D})\approx\ell({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}},\mathcal{D})+\underbrace{\nabla\ell({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}})^{\top}}_{=0}({{\boldsymbol{\mathbf{\theta}}}}-{{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}})+\frac{1}{2}({{\boldsymbol{\mathbf{\theta}}}}-{{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}})^{\top}\nabla^{2}\ell({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}})({{\boldsymbol{\mathbf{\theta}}}}-{{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}).$ (42)
Since ${{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}$ is a local maximum, the gradient term vanishes. Identifying the remaining quadratic form with the log-density of a Gaussian $\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}},{{\boldsymbol{\mathbf{\Sigma}}}})$ yields the approximation parameters:
$q({{\boldsymbol{\mathbf{\theta}}}})$ $=\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}},{{\boldsymbol{\mathbf{\Sigma}}}}),$ (43)
${{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}$ $=\arg\max_{{{\boldsymbol{\mathbf{\theta}}}}}\ell({{\boldsymbol{\mathbf{\theta}}}},\mathcal{D}),$ (44)
${{\boldsymbol{\mathbf{\Sigma}}}}$ $=-\left[\nabla^{2}_{{{\boldsymbol{\mathbf{\theta}}}}{{\boldsymbol{\mathbf{\theta}}}}}\ell({{\boldsymbol{\mathbf{\theta}}}},\mathcal{D})\Big\rvert_{{{\boldsymbol{\mathbf{\theta}}}}={{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}}\right]^{-1},$ (45)
where the covariance $\boldsymbol{\mathbf{\Sigma}}$}} is defined as the negative inverse Hessian.
However, computing the full Hessian is unfeasible in practice. The Generalized Gauss-Newton (ggn) approximation is commonly adopted for tractability. Under the GGN, the per-data-point Hessian is approximated using the Fisher Information:
$\nabla^{2}_{{{\boldsymbol{\mathbf{\theta}}}}{{\boldsymbol{\mathbf{\theta}}}}}\log p(y\mid{{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}))\approx-J_{{\boldsymbol{\mathbf{\theta}}}}({{\boldsymbol{\mathbf{x}}}})^{\top}{{\boldsymbol{\mathbf{\Lambda}}}}({{\boldsymbol{\mathbf{f}}}})J_{{\boldsymbol{\mathbf{\theta}}}}({{\boldsymbol{\mathbf{x}}}}),$ (46)
where ${{\boldsymbol{\mathbf{\Lambda}}}}({{\boldsymbol{\mathbf{f}}}})=-\mathbb{E}_{y\sim p(y|{{\boldsymbol{\mathbf{f}}}})}[\nabla_{ff}^{2}\log p(y|{{\boldsymbol{\mathbf{f}}}})]\in\mathbb{R}^{C\times C}$ is the Hessian of the negative log-likelihood w.r.t the logits (which coincides with the Fisher information metric for canonical link functions), and $J_{{\boldsymbol{\mathbf{\theta}}}}({{\boldsymbol{\mathbf{x}}}})\in\mathbb{R}^{C\times P}$ is the Jacobian of the outputs w.r.t the parameters.
Following Immer et al. (2021), the ggn approximation enables a local linearization of the network $f_{{\boldsymbol{\mathbf{\theta}}}}({{\boldsymbol{\mathbf{x}}}})$ around $\boldsymbol{\mathbf{\theta}}$$^{*}$ (usually taken as ${{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}}$):
${{\boldsymbol{\mathbf{f}}}}_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}})={{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}^{*})+J_{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}})({{\boldsymbol{\mathbf{\theta}}}}-{{\boldsymbol{\mathbf{\theta}}}}^{*}).$ (47)
This reduces the bnn to a multi-output Generalized linear model (glm). Comparing the two predictive distributions:
$p_{\text{BNN}}(y\mid{{\boldsymbol{\mathbf{x}}}},\mathcal{D})$ $=\mathbb{E}_{q({{\boldsymbol{\mathbf{\theta}}}})}\Big[p(y\mid{{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}))\Big],$ (48)
$p_{\text{GLM}}(y\mid{{\boldsymbol{\mathbf{x}}}},\mathcal{D})$ $=\mathbb{E}_{q({{\boldsymbol{\mathbf{\theta}}}})}\Big[p(y\mid{{\boldsymbol{\mathbf{f}}}}_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}))\Big].$ (49)
Notably, $p_{\text{GLM}}$ pushes the posterior Gaussian through the linearized function, resulting in a Gaussian distribution in logit space, while $p_{\text{BNN}}$ uses the original non-linear network.
B.2 Uncertainty quantification
To decompose the uncertainty we use the law of total variance on the one-hot encoded target variable $\boldsymbol{\mathbf{y}}$$^{*}$. Considering a test point $\boldsymbol{\mathbf{x}}$$^{*}$, the predictive covariance (tu) is:
[…]
where $\mathbb{E}[{{\boldsymbol{\mathbf{y}}}}^{*}\mid{{\boldsymbol{\mathbf{x}}}}^{*},{{\boldsymbol{\mathbf{\theta}}}}]=\text{softmax}({{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}}^{*};{{\boldsymbol{\mathbf{\theta}}}}))$ corresponds to the model’s output probabilities. Both au and eu are $C\times C$ positive semi-definite matrices.
  • The au matrix captures the expected inherent noise (irreducible aleatoric uncertainty).
  • The eu matrix captures the model’s uncertainty regarding the true class probabilities (epistemic uncertainty).
To make the computation explicit, we assume an isotropic Gaussian prior $p({{\boldsymbol{\mathbf{\theta}}}})=\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}};\mathbf{0},\lambda^{-1}{{\boldsymbol{\mathbf{I}}}})$, implying $\nabla^{2}_{{{\boldsymbol{\mathbf{\theta}}}}{{\boldsymbol{\mathbf{\theta}}}}}\log p({{\boldsymbol{\mathbf{\theta}}}})=-\lambda{{\boldsymbol{\mathbf{I}}}}$. Consequently, the posterior covariance $\boldsymbol{\mathbf{\Sigma}}$}} is determined by the sum of the data-induced curvature (Fisher information) and the prior precision:
${{\boldsymbol{\mathbf{\Sigma}}}}=\left(\sum_{n=1}^{N}J_{{{\boldsymbol{\mathbf{x}}}}_{n}}^{\top}{{\boldsymbol{\mathbf{\Lambda}}}}({{\boldsymbol{\mathbf{f}}}}_{n})J_{{{\boldsymbol{\mathbf{x}}}}_{n}}+\lambda{{\boldsymbol{\mathbf{I}}}}\right)^{-1}.$ (51)
We study the predictive covariance in logit space, which serves as a proxy for the eu in probability space of Equation˜50. Using the linearized predictive $p_{\text{GLM}}$, the covariance matrix is given by:
$\mathbf{\Sigma}_{\text{logit}}({{\boldsymbol{\mathbf{x}}}})$ $=\mathbb{V}_{q({{\boldsymbol{\mathbf{\theta}}}})}[{{\boldsymbol{\mathbf{f}}}}_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}})]$ (52)
$=\mathbb{V}\Big[{{\boldsymbol{\mathbf{f}}}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}^{*})+J_{{\boldsymbol{\mathbf{x}}}}({{\boldsymbol{\mathbf{\theta}}}}-{{\boldsymbol{\mathbf{\theta}}}}^{*})\Big]$ (53)
$=J_{{\boldsymbol{\mathbf{x}}}}\mathbb{V}[{{\boldsymbol{\mathbf{\theta}}}}]J_{{\boldsymbol{\mathbf{x}}}}^{\top}$ (54)
$=J_{{\boldsymbol{\mathbf{x}}}}{{\boldsymbol{\mathbf{\Sigma}}}}J_{{{\boldsymbol{\mathbf{x}}}}}^{\top}$ (55)
$=J_{{\boldsymbol{\mathbf{x}}}}\left(\sum_{n=1}^{N}J_{{{\boldsymbol{\mathbf{x}}}}_{n}}^{\top}{{\boldsymbol{\mathbf{\Lambda}}}}({{\boldsymbol{\mathbf{f}}}}_{n})J_{{{\boldsymbol{\mathbf{x}}}}_{n}}+\lambda{{\boldsymbol{\mathbf{I}}}}\right)^{-1}J_{{\boldsymbol{\mathbf{x}}}}^{\top}\in\mathbb{R}^{C\times C}.$ (56)
The scalar epistemic uncertainty is obtained by taking the trace of this covariance matrix (summing the variances across output dimensions):
$\text{EU}_{\text{logit}}({{\boldsymbol{\mathbf{x}}}})=\text{Tr}\left(\mathbf{\Sigma}_{\text{logit}}({{\boldsymbol{\mathbf{x}}}})\right).$ (57)
The term $J_{{{\boldsymbol{\mathbf{x}}}}_{n}}^{\top}{{\boldsymbol{\mathbf{\Lambda}}}}({{\boldsymbol{\mathbf{f}}}}_{n})J_{{{\boldsymbol{\mathbf{x}}}}_{n}}$ represents the precision contributed by the training data. If the test Jacobian $J_{{\boldsymbol{\mathbf{x}}}}$ is orthogonal to the subspace spanned by the training gradients, the covariance reduces to the prior term $\lambda^{-1}J_{{\boldsymbol{\mathbf{x}}}}J_{{\boldsymbol{\mathbf{x}}}}^{\top}$.
As highlighted by Immer et al. (2021), the Bayesian glm in weight space is equivalent to a gp in function space with a particular kernel.
B.3 Experiments: shallow models
Figure 12: Scaling Analysis of lla: (Left) Growth of the Hessian (data precision) spectrum as a function of dataset size $N$. Both the maximum eigenvalue (blue) and the average eigenvalue (cyan) scale linearly with $N$, confirming that information accumulates additively. (Right) Average eu on the test set for varying prior precisions $\lambda\in[0.001,1]$. The dashed (gray) line indicates the theoretical $O(N^{-1})$ reference scaling. The legend displays the empirical scaling exponent $\alpha$ for each prior setting (where $\text{EU}\propto N^{-\alpha}$), showing that all curves asymptotically converge to the data-driven $N^{-1}$ regime regardless of the prior strength.
Dataset. We use the Two Moons dataset, a binary classification problem ($C=2$) with inputs ${{\boldsymbol{\mathbf{x}}}}\in\mathbb{R}^{2}$. To observe the transition from the prior-dominated to the data-dominated regime, we generated training sets of varying sizes $N\in\{5,10,20,50,100,200,500\}$. A fixed independent test set of $N_{\text{test}}=100$ samples was used for evaluation.
Model Architecture. We employed an mlp with a single hidden layer and $\tanh$ activation functions. The default architecture consisted of $50$ hidden units. The model parameters map were estimated via standard regularized cross-entropy minimization using the Adam optimizer. All experiments were repeated over 3 independent folds with different random seeds for data generation and weight initialization.
We computed the uncertainty using the Linearized Laplace approximation (GLM predictive):
  1. 1.
    Hessian Approximation: We use the ggn approximation to the Hessian and computed the full Hessian covariance matrix ${{\boldsymbol{\mathbf{\Sigma}}}}\in\mathbb{R}^{P\times P}$ (where $P$ is the number of parameters).
  2. 2.
    Linearization: The predictive distribution was treated as a glm by linearizing the network function around the map estimate.
  3. 3.
    Prior: We train models with different prior’s precision $\lambda$ to observe the shift in the "knee" of the scaling curve (the transition from prior to data dominance).
  4. 4.
    Uncertainty Metrics: We quantify eu using two complementary metrics to capture uncertainty in both function and probability space:
    • Variance (logit space): For a given test point $\boldsymbol{\mathbf{x}}$}}, we compute the trace of the predictive covariance in logit space. This serves as a proxy for the volume of the posterior in function space:
      $\text{EU}_{\text{logit}}({{\boldsymbol{\mathbf{x}}}})$ $=\mathbb{V}_{q({{\boldsymbol{\mathbf{\theta}}}})}[{{\boldsymbol{\mathbf{f}}}}_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}})]$ (58)
      $=\text{Tr}\left(J_{{\boldsymbol{\mathbf{x}}}}{{\boldsymbol{\mathbf{\Sigma}}}}J_{{\boldsymbol{\mathbf{x}}}}^{\top}\right),$ (59)
      where $J_{{\boldsymbol{\mathbf{x}}}}\in\mathbb{R}^{C\times P}$ is the Jacobian of the model output w.r.t parameters.
      Theoretical expectations: Assuming the training data is i.i.d., the accumulated Fisher information grows linearly with NN. Consequently, in the data-dominated regime (large NN), the posterior variance scales as ${{\boldsymbol{\mathbf{\Sigma}}}}_{N}\propto\frac{1}{N}$. See Figure˜12.
    • Variance (probability space): In the binary setting, this metric captures the variance of the predicted Bernoulli parameter $\pi=\sigma(f)$ induced by the posterior on the weights. Since the mapping from logit space to probability space is non-linear, a Gaussian distribution on $f$ does not result in a Gaussian on $\pi$. The variance is given by:
      $\text{EU}_{\text{var}}({{\boldsymbol{\mathbf{x}}}})=\mathbb{V}_{q({{\boldsymbol{\mathbf{\theta}}}})}\Big[\sigma\big(f_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}})\big)\Big].$ (60)
      Unlike the logit-space variance, this quantity is bounded in $[0,0.25]$. Because the sigmoid saturates, high variance in logit space does not necessarily translate to high variance in probability space if the mean is far from the decision boundary. This expectation is typically intractable and is approximated via Monte Carlo sampling using $S$ samples from $q({{\boldsymbol{\mathbf{\theta}}}})$:
      $\text{EU}_{\text{var}}({{\boldsymbol{\mathbf{x}}}})\approx\frac{1}{S-1}\sum_{s=1}^{S}\left(\sigma(f^{(s)})-\bar{\pi}\right)^{2},\quad\text{where }f^{(s)}\sim q({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}},{{\boldsymbol{\mathbf{\Sigma}}}}).$ (61)
      Theoretical expectations: Applying the Delta method (first-order Taylor expansion) to the sigmoid function, the variance in probability space is approximated by: $\mathbb{V}[\sigma(f)]\approx\left[\frac{\partial\sigma(f)}{\partial f}\right]^{2}\mathbb{V}[f]=\Big[\pi(1-\pi)\Big]^{2}\cdot\text{EU}_{\text{logit}}({{\boldsymbol{\mathbf{x}}}}).$ (62) While the asymptotic rate remains $O(N^{-1})$, the magnitude is heavily dampened by the term $\pi(1-\pi)]^{2}$. As the model becomes more confident (predictions $\pi$ close to 0 or 11), the gradient of the sigmoid vanishes, suppressing the visible eu even if the underlying logit uncertainty is moderate.
    • Entropy (probability space): We also measure the eu in probability space via the Mutual Information (mi) between predictions and parameters. This is estimated using Monte Carlo sampling with $S$ samples from the glm predictive:
      $\text{EU}_{\text{ent}}({{\boldsymbol{\mathbf{x}}}})=H\left[\frac{1}{S}\sum_{s=1}^{S}{{\boldsymbol{\mathbf{p}}}}_{s}\right]-\frac{1}{S}\sum_{s=1}^{S}H[{{\boldsymbol{\mathbf{p}}}}_{s}],$ (63)
      where $H[\cdot]$ denotes the Shannon entropy, ${{\boldsymbol{\mathbf{p}}}}_{s}=\text{softmax}({{\boldsymbol{\mathbf{f}}}}_{\text{lin}}^{{{\boldsymbol{\mathbf{\theta}}}}^{*}}({{\boldsymbol{\mathbf{x}}}};{{\boldsymbol{\mathbf{\theta}}}}_{s}))$, and ${{\boldsymbol{\mathbf{\theta}}}}_{s}\sim\mathcal{N}({{\boldsymbol{\mathbf{\theta}}}}_{\text{MAP}},{{\boldsymbol{\mathbf{\Sigma}}}})$.
      Theoretical expectations: Similar to the variance metric, the mi scales asymptotically as $O(N^{-1})$. By applying a second-order Taylor expansion to the entropy functional around the mean prediction $\bar{{{\boldsymbol{\mathbf{p}}}}}$, the mi can be approximated as: $\text{EU}_{\text{ent}}({{\boldsymbol{\mathbf{x}}}})\approx\frac{1}{2}\text{Tr}\left(\mathcal{I}(\bar{{{\boldsymbol{\mathbf{p}}}}})\cdot\mathbb{V}[{{\boldsymbol{\mathbf{p}}}}]\right),$ (64) where $\mathcal{I}(\bar{{{\boldsymbol{\mathbf{p}}}}})$ is the Fisher Information of the observation model evaluated at the mean prediction. Since the predictive variance $\mathbb{V}[{{\boldsymbol{\mathbf{p}}}}]$ is driven by the posterior weight variance ${{\boldsymbol{\mathbf{\Sigma}}}}\propto N^{-1}$, the entropy-based uncertainty inherits this scaling, converging linearly to zero in the large data limit.
Appendix C Additional results
@@
Figure 13: Vision Transformer (ViT) on ImageNet-32 dataset: We use mcd with fixed dropout rate $p=0.1$. We train for $200$ epochs using sgd with cosine annealing. In (a) we report uncertainty scaling on subsets from size $250K$ up to $1.2M$. In (b) we discard the first point highlighting the decreasing rate of eu for large values of $N$.
@@
Figure 14: WideResNets ($\mathbf{w}$$\mathbf{d}$) on CIFAR-100 dataset: We train the models with mcd with fixed dropout rate $p=0.2$ (first row) and $p=0.5$ (second row). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data and we show the mean uncertainty for each NN subset. Dashed lines represent linear regression fitted to the test mean of each uncertainty metric. Both axes are on a logarithmic scale.
@@
Figure 15: WideResNets ($\mathbf{w}$$\mathbf{d}$) on CIFAR-10 dataset - extrapolation: We report uncertainty trends for different WideResNets and dropout rates. We mention in the main paper that the scaling laws we derive in this context are practically useful to extrapolate uncertainties to NN arbitrarily large. We report $\mathbb{V}[\sigma(\mathbf{f})]$, the variance computed over the Softmax predictions of the mc samples, averaged over all the test samples ($\times$ in the plots). The black dashed lines have slope $\gamma_{f}$, the blue dashed lines correspond to two thresholds of predictive uncertainty.
@@
Figure 16: ResNets on CIFAR-100 dataset with de: We train the models with de with ensemble members $M=5$ (first row) and $M=10$ (second row). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data and we show the mean uncertainty for each NN subset. Dashed lines represent linear regression fitted to the test mean of each uncertainty metric. Both axes are shown on a logarithmic scale.
@@
Appendix D Experimental setup
Table 1: Vision Transformer (ViT) architecture: To obtain the results in Figure˜8 and Figure˜13, we use a compact ViT model with details specified in the table below. The implementation is available at https://github.com/kentaroy47/vision-transformers-cifar10.